The probability distribution for the number of automobiles lined up at a Lakeside Olds dealer at opening time (7:30 a.m.) for service is:
 
Number 1 2 3 4
Probability 0.05 0 .30 0.40 0.25
 
On a typical day, how many automobiles should Lakeside Olds expect to be lined up at opening time?
Multiple Choice
  • 10.00
  • 1.00
  • 2.85
Answer #1

Solution :

ecpect value = = X * P(X)

= 1 *0.05 + 2 * 0.30+ 3 *0.40 + 4 * 0.25

= ( 0.05+ 0.60 + 1.2 + 1)

= 2.85

correct option 2.85